datonekidintheclass:

Which is the equation of a trend line that passes through the points (7, 450) and (14, 401)? y = negative 7 x + 499 y = negative StartFraction 1 Over 7 EndFraction x + 451 y = StartFraction 1 Over 7 EndFraction x + 449 y = 7 x + 401

jhonyy9:

do you know what mean that a trend line passes through the points (7, 450) and (14, 401)? what are these points for this equation ?

ramen:

do you know what trend means?

Mercury:

apply the slope formula for the points (7, 450) and (14, 401), then plug the slope into the point-slope formula. re-arrange into y = mx + b form. slope formula: (y2 - y1) / (x2 - x1) where (x1,y1) and (x2,y2) are your points. it doesn't matter which one you asisgn to be x1,y1 and which one you assign to be x2,y2 but be consistent with your choice. point slope formula: y - y1 = m(x-x1) where (x1,y1) is simply one of the points